HUBBLE SPACE TELESCOPE DAILY REPORT #1262 PERIOD COVERED: 0000Z Daily Status Report as of 313/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 5560 (WF/PC-2 Cycle 4 Calibration: Internal Monitor. 1) The WF/PC-2 was used to make several internal bias, K-Spot, and flat calibration observations to monitor the CCD camera electronics, the optical alignment, and to check for contamination. No other anomalies were noted. 1.2 Completed FOS/RD 5664 (Quasar Absorption Line Survey: Cycle 4 High) The Faint Object Spectrograph (Red Side) was used to make spectrographic observations of quasi-stellar object UM425 to look for absorption lines due to gas in the intergalactic medium and galactic halos. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed WF/PC-2 5505 (Deep Probes of Gravitational Lens Systems) The WF/PC-2 was used to make deep images of of the elliptical galaxy MG0414+053 which is a member of the Coma Cluster. The intent is to improve our understanding of the stellar populations in ellipticals and to determine the distance to the Coma Cluster. The observations were completed and planned, and no problems were noted. 1.4 Completed WF/PC-2 5764 (WF/PC-2 Cycle 4 Calibration: Internal Flats - Repeat) The WF/PC-2 was used to make internal flat calibration images using the VIS lamp and internal sources. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed WF/PC-2 5571 (WF/PC-2 Cycle 4 Calibration: Earthflats, Smaller Set) The WF/PC-2 was used to make several ultraviolet and narrow-band filter Earth flat calibration observations to correct for the telescope's illumination pattern. The observations were completed as planned, and no problems were reported. 1.6 Completed WF/PC-2 5351 (Properties of Intermediate Redshift Galactic Disks Causing Damped Lyman-Alpha) The WF/PC-2 was used to make observations to study the diffuse material at intermediate redshifts that is identified by absorbing Lyman-Alpha spectrum from quasar 3C286.0. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 5 Successful : 5 Scheduled Reacquisitions : 9 Successful : 9 2.2 FHST Updates: Scheduled : 11 Successful : 11 2.3 Significant HSTARs: None 2.4 Operations Notes: "H" telemetry format check-out activities continued in the background throughout the day. 3.0 ANOMALY STATUS: 3.1 Significant New Anomalies: None 3.2 Significant Unresolved Anomalies: None 4.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-